OpenStudy (anonymous):

Part 1: Solve the inequality −5x + 8 less than or equal to 5x − 2 and show all of your work. You can use <= to represent the less than or equal to symbol. Part 2: Use complete sentences to describe the graph of the solution from Part 1. Be sure to include the type of circle and which direction the shading will be.

OpenStudy (anonymous):

someone please help

OpenStudy (anonymous):

-5x+8=5x-2 +5x +5x ______________ 8 = 10x-2 -2 -2 ------------- 10 = 10x --- --- 10 10 x=1 Check your Work -5(1)+8=3 5(1)-2=3